I have to figure out how much no fat yogurt should be mixed with regular yogurt to obtain 60 pounds of low fat yogurt.
I have 3% fat yogurt blends with no-fat yogurt that yields a 1%low fat yogurt.
I know that x60 is what I need to determine.
.03x is the 3%, .01 is the 1% which gives me 4% total lowfat yogurt. Then I divided that into the 60 needed and I came up with 17 pounds.

Answer by richwmiller(17219)

Not quite
We want to end up with 60 pounds of 1%
The no fat will have 0% fat that will be x
y will be how much of the 3% we add
x+y equals 60
.03*y=.01*(60)
x+y=60
y=20 x=40
check
3/100*20+40*0%=1/100*60
60/100+0=60/100
.6=.6
ok
another way to look at it is from the non fat view
1=100%
.97*y+1*x=60*.99
x+y=60
same answers
y=20 x=40
40 pounds of non fat